evaluate (a + b).(a - b) if a = 2i - 3j + 5k and b = 3i + j - 2k
Answer by Edwin McCravy(20055)

(a + b)•(a - b) if a = 2i - 3j + 5k and b = 3i + j - 2k

a + b = (2i - 3j + 5k) + (3i + j - 2k) = 5i - 2j + 3k

a - b = (2i - 3j + 5k) - (3i + j - 2k) = -i - 4j + 7k

(a + b)•(a - b) = 

(5i - 2j + 3k)•(-i - 4j + 7k) = (5)(-1)+(-2)(-4)+(3)(7) =

24

Edwin